The perimeter of a rectangular car floor mat is 190cm. The width is a 1/4 of length. Find the length and width of the table.
Found 2 solutions by Boreal, Alan3354:
Answer by Boreal(15235)

width=x
length =4x
2W+2L=perimeter=190 cm
2x+8x=190
10x=190 cm
x=19 cm
it is 19 cm wide and 76 cm long

Answer by Alan3354(69443)

The perimeter of a rectangular car floor mat is 190cm. The width is a 1/4 of length. Find the length and width of the table.
--------------
2W + 2L = 190
L = 4W
----
2W + 8W = 190
etc.
==================
What table?